Suppose that you are the Manager of a soccer stadium and you are


responsible for the sale of tickets. Two matches are scheduled to be


played during the next fortnight, the first between team X and team Y.


Market research indicates that you can sell 40 000 tickets for the X-Y


clash at R10 each, or 30 000 tickets at R20 each. Which option would


you choose? What is the arc price elasticity of the demand for tickets for


this particular game?


1
Expert's answer
2021-03-29T12:17:45-0400



(a) 30 000 tickets at R20 each


Ed ="\frac{40000 - 30000}{10-20}" "\times" "\frac{10+20}{40000+30000}"

"{-1000}\times""\frac{30}{70000}" ="\frac{-3}{7}"

Therefore, the demand is inelastic.